OpenStudy (anonymous):

Suppose that a family wants to start a college fund for their child. If you can get an APR of 7.5% and want the fund to have a value of $75,000 after 18 years, how much should you deposit monthly? Assume an ordinary annuity.

OpenStudy (amistre64):

you basically want to loan yourself money ... how would you calculate the monthly payments if you took out a loan with that information?

OpenStudy (amistre64):

Bn = Bo k^n - P(1-k^n)/(1-k) assuming we want a balance of 0 after n periods 0 = Bo k^n - P(1-k^n)/(1-k), and solving for P Bo k^n = P(1-k^n)/(1-k) Bo k^n(1-k)/(1-k^n) = P

OpenStudy (amistre64):

k is the compounding interest stuff so in this case its (1+.075/12) and n=18*12

OpenStudy (amistre64):

so, 75000k^(18*12)(1-k)/(1-k^(18*12)), k=(1+.075/12) http://www.wolframalpha.com/input/?i=75000k%5E%2818*12%29%281-k%29%2F%281-k%5E%2818*12%29%29%2C+k%3D%281%2B.075%2F12%29 about 634 id say

OpenStudy (amistre64):

of course your class might want you to use their formulas and tables ... but i never liked them :)

OpenStudy (anonymous):

Suppose that a family wants to start a college fund for their child. If you can get an APR of 7.5% and want the fund to have a value of $75,000 after 18 years, how much should you deposit monthly? Assume an ordinary annuity. a. $164.98 c. $166.21 b. $165.30 d. $167.52

OpenStudy (anonymous):

This class is confusing! None of this makes any sense!

OpenStudy (amistre64):

hmm, let me review my thoughts and see if its still correct

OpenStudy (amistre64):

.075 is good, 12 months a year .... for 18 years

OpenStudy (amistre64):

1 + 2 + 3 + ... + n-1 -2 -3 - ... - n-1 -n --------------------- 1-k^n is fine

OpenStudy (amistre64):

75000 = P(1-k)/(1-k^n) thats better, 164.98

OpenStudy (amistre64):

i was compounding a balance that did not exist instead of working up to it i spose. we dont have the 75000 compounding since we didnt actually take a loan out.

OpenStudy (anonymous):

Thank you! (:

OpenStudy (amistre64):

sure :)